OpenStudy (anonymous):

A tank can be filled by one pipe in 45 minutes and by another in 55 minutes. How long will it take both to fill it?

OpenStudy (anonymous):

does x = time required for both pipes to fill tank?

OpenStudy (anonymous):

x = 24 1/3

OpenStudy (anonymous):

thanks for the help MSFFL

OpenStudy (anonymous):

I'd give you a badge but it isn't letting me! Which is strange

OpenStudy (anonymous):

1/t = 1/55 + 1/45 = (9 + 11)/ 495 => t = 495/ 20 = 24. 75 = 24min 45 sec